OpenStudy (anonymous):

simplify (-9x)97p)+(-3x)(-5p)

OpenStudy (anonymous):

* (7p) it should be

OpenStudy (austinl):

That is not what I got, sorry. We have to multiplications of two variables that are then added together. That means that there are going to be two variables in the answer. \[((-9x) \times 97p)+(-3x \times (-5p))\]

OpenStudy (anonymous):

no in the equation it should not be 97p) it should be (7p)

OpenStudy (anonymous):

i was correcting my mistake lol

OpenStudy (austinl):

\[((-9x) \times 7p) + ((-3x) \times (-5p))\] So what is -9 times 7? And, -3 times -5?

OpenStudy (anonymous):

i dont ahve a calcualtor on me atm

OpenStudy (anonymous):

could you help

OpenStudy (austinl):

sorry, i got distracted. -9*7=-63 -5*-3=15 so we have -63xp + 15xp and that gives us -48xp

OpenStudy (anonymous):

thanks lol i got distracted too

OpenStudy (austinl):

coolio, glad i could help!

OpenStudy (anonymous):

would you help me on my next questions please